Bear population. To estimate the size of the bear population on the Keweenaw Peninsula, conservationists captured, tagged, and released 50 bears. One year later, a random sample of 100 bears included only 2 tagged bears. What is the conservationist's estimate of the size of the bear population?
To estimate the size of the bear population on the Keweenaw Peninsula, conservationists captured, tagged, and released 50 bears. One year later, a random sample of 100 bears included only 2 tagged bears. What is the conservationist's estimate of the size of the bear population? Let "x" be the number in the population. Real population / sample caught = released population / sample caught.
x/100=50/2
x=2500 bears
Tagged by conservationists = 50 bears which represented an estimated 2% of the population = 50/.02=2500 total bear population.







Page 437 problem 56 and Page 444 problem 10
Problem: 10 pg. 444
For the second problem in this assignment I am asked to solve this equation for y. The first thing I notice is that it is a single fraction (ratio) on both sides of the equal sign so basically it is a proportion which can be solved by cross multiplying the extremes and means.

This is the proportion set up and ready to solve. I will cross multiply the x and y setting the extremes equal to the means.
(y-1)*4=(x+3)-(-3)
4y-4=-3x-9
4y=-3x-9+4
4y=-3x-5 Divide both sides by 4
4y = -3x – 5 Simplify
4 4 4
4y = -(3) * x – 5
4 (4) 4
y = -(3) * x-5 Answer to problem 10 on page 444.
(4) 4


IS THIS CORRECT OF THE WORK i PERFORMED SO FAR?

Answer by stanbon(75887)

This is the proportion set up and ready to solve. I will cross multiply the x and y setting the extremes equal to the means.
(y-1)*4=(x+3)(-3)
4y-4=-3x-9
4y=-3x-9+4
4y=-3x-5 Divide both sides by 4
------
y = (-3/4)x - (5/4)
======================
Corrections made.
=================
Cheers,
Stan H.
=======================